OpenStudy (anonymous):

Write the standard form of the equation of the line that passes through the point (-1, 3) and is parallel to the graph of 2x - 7y + 1 = 0.

OpenStudy (anonymous):

take 2x-7y+1=0 and put it into slope intercept form(y=mx+b) 2x+1=7y (2x+1)/7=y y=(2/7)x+1/7, the slope would be 2/7 since it is parallel then you put it into point slope form y-y1=m(x-x1) y-3=2/7(x-(-1)) y-3=(2/7)x+2/7 y=(2/7)x+ 3 2/7 standard form is Ax+By+C=0 0=(2/7)x-y+3 2/7